OpenStudy (anonymous):

what are the roots of the equation (x^3+4x^2+6x+12)=0 how am i supposed to find this ?? help....pleaseeee

hartnn (hartnn):

any similar type of sum done before ?

OpenStudy (anonymous):

do u know about hit and trail method

hartnn (hartnn):

roots of this are horrible....can't apply hit and trail... is the question correct ?@erica.d

OpenStudy (anonymous):

no this is the first one...and i haven't done this before...i don't know about hit and trial :(

OpenStudy (anonymous):

do i go for calculator?

OpenStudy (hba):

For a cubic equation like: Ax^3 + Bx^2 + Cx + D = 0, having three roots as a, b, c; the relations between these roots are given by: a*b*c = -D/A a+b+c = -B/A ab+bc+ca = C/A

OpenStudy (anonymous):

that's horrible

OpenStudy (hba):

Have you ever tried solving Quad roots ?

OpenStudy (anonymous):

yes...but this is cubic

OpenStudy (hba):

I know I jus Wanted to Know Because Solving Cubic Roots Is Diff

OpenStudy (anonymous):

@satellite73 .....

OpenStudy (hba):

So, for the equation, x^3+4x^2+6x+12=0 A=1,B=4,C=6,D=12 The product of the roots is: abc= -D/A = -12/1=-12

OpenStudy (hba):

The sum of the roots is : a+b+c= -B/A=-4/1=-4

hartnn (hartnn):

i would like to see how @satellite73 approaches this............

OpenStudy (anonymous):

@satellite73 pleaseeeeee @satellite73 and thank you @hba

OpenStudy (anonymous):

@cwrw238

hartnn (hartnn):

@ganeshie8

OpenStudy (hba):

I am Pretty Sure I am correct :)

OpenStudy (anonymous):

i guess i'll do it in a while :)

hartnn (hartnn):

@mukushla

OpenStudy (anonymous):

analyticall approach will be painfull

OpenStudy (hba):

LoL @mukushla

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

hba ur method also will be back to original equation...

OpenStudy (anonymous):

why you guys are laughing ?? :(

OpenStudy (anonymous):

nothing we can change the equation to the form\[x^3+px+q=0\]and this one is solved by hartnn where is ur toturial hartnn?

OpenStudy (hba):

I know that :)

hartnn (hartnn):

but how do we change ?

OpenStudy (anonymous):

for \[ax^3+bx^2+cx+d=0\]\[t=x+\frac{b}{3a}\]will change the equation to\[x^3+px+q=0\]if im not wrong

OpenStudy (anonymous):

but numerical methods like newton-raphson is better for such equations

OpenStudy (nipunmalhotra93):

here's is a method... find the derivative of the function. You can see that is always positive. So, the function is monotonically increasing. So it will have only 1 real root. The other two roots are complex. So the roots are r, p+iq and p-iq Put these in the equations given by @hba and you have 3 equations and 3 variables. Solve for r,p and q.

OpenStudy (anonymous):

here is a nice article http://en.wikipedia.org/wiki/Cubic_function

hartnn (hartnn):

as already told, maybe solving those will give back u original question...

OpenStudy (nipunmalhotra93):

let me check...

OpenStudy (nipunmalhotra93):

hmmm... you're right :\

OpenStudy (anonymous):

i think when we encounter cubic equations with integer coefficients better to go with rational root theorem directly...and checking for possible rational roots. if there is not any id say just using numerical methods

OpenStudy (anonymous):

otherwise if u want solve it analyticall it will be painfull :)